The stock of Bell Towers has an expected return of 14.7%. The expected market risk premium is 7% and the risk-free rate is 3.8%. What is the beta of Bell Towers' stock? A. 1.23 B. 1.33 C. .92 D. 1.56 E. 1.41
Answer #1

D. 1.56

Expected return = Risk-free rate + Beta(Market risk premium)

0.147 = 0.038 + Beta(0.07)

0.147 - 0.038 = Beta(0.07)

Beta(0.07) = 0.109

Beta = 0.109 / 0.07

Beta = 1.56